OpenStudy (anonymous):

Solve using addition principal 7.2=s-2.5

OpenStudy (mertsj):

Add 2.5 to both sides.

OpenStudy (anonymous):

The addition principle states that if a=b, then a+c=b+c 7.2+2.5=s-2.5+2.5 9.7=s